OpenStudy (anonymous):

Need some help with negative exponents please! Simplify. x^0yz^-3 Simplify. 6x^-2 Simplify. f^-1g^-1 Simplify. y^-3 Simplify. (-2)^-3 I got these wrong on one of my assignments :[

OpenStudy (jdoe0001):

just keep in mind that \(\bf a^{-{\color{red} n}} = \cfrac{1}{a^{\color{red} n}}\) and expand accordingly

OpenStudy (anonymous):

what are those red things? @jdoe0001

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (jdoe0001):

\(\bf x^0yz^{-{\color{red}{ 3}}}\implies x^0\cdot y\cdot \cfrac{1}{z^{\color{red}{ 3}}}\implies 1\cdot y\cdot \cfrac{1}{z^{\color{red}{ 3}}}\implies y\cdot \cfrac{1}{z^{\color{red}{ 3}}} \\ \quad \\ \cfrac{y}{z^{\color{red}{ 3}}}\) so... you'd convert the negative exponent like so, and then simplify :)

OpenStudy (anonymous):

thank you! @jdoe0001

OpenStudy (jdoe0001):

yw